Hello

i need help for this!Let A be the matrix below. Find the singular values of A and enter them as a comma-separated list. Use these to find a singul
Answer #1

Given that A = ri 17 -2 -2 AT = 11 -2 -17 Now find Ant pi 1751-2-17 -2-2 1-2 U rixlfix1 16271(-2) 16-1)+1X1 (-2)x1+12)/ (-2)(- (0-1) [16-81-81+12) + 4[87417 - 38-168 - 48+212-166 +872 +212+(-43)+ (-32+144] = -13 +1212-201=0. [-12 +124–20] =0 1 = -12+